OpenStudy (anonymous):

determine whether the function is even, odd, or neither f(X)=-X^3+4X^2

OpenStudy (zzr0ck3r):

if its even \(f(-x) = f(x)\) and here \(f(-x) = x^3+4x^2\) so it is NOT even if it is odd \(f(-x) = -f(x)\) and \(f(-x) = x^3+4x^2\) is this the same thing as \(-f(x) = -(-x^3+4x^2)=x^3-4x^2\)?

OpenStudy (anonymous):

im not sure im alittle confused